The freezing point of benzene is 5.5oC. What is the freezing point of a solution of 5.00 g of naphthalene (C10H8) in 444 g of benzene (Kf of benzene = 4.90oC/m


1
Expert's answer
2022-02-07T23:07:01-0500

Mass of benzene in Kg=444÷1000

=0.444kg

Moles of naphthalene= mass/molar mass

=5/128= 0.039moles.

M=moles of naphthalene/mass of benzene

M=0.039÷0.444=0.088mol/kg

Vant Hoff factor of naphthalene=1

Tf= iKfM

Tf= 1× 4.9×0.088

=0.43oC